OpenStudy (anonymous):

PLEASE HELP Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (anonymous):

hi use quadratic formula in those two equation, do you know thw quadratic formula?

OpenStudy (anonymous):

X^2 + 5x + 6 = 0 this can be factorised as X^2 + 2x + 3x + 6 = 0 X(x+2) + 3 ( x + 2) = 0 (x+2) * ( x + 3 ) = 0 x = -2 or -3 For x^2 + 4x + 4, it is factorised as x^2 + 2x + 2x + 4 =0 this simplifies to (x+ 2) * (x+2) =0 or (x+2)^2 =0 in both cases, x = -2

OpenStudy (anonymous):

Thank you so much for the help

OpenStudy (anonymous):

Thanks Mark for trying to help but i guess ishan beat u too it lol

OpenStudy (anonymous):

try also learning to use the quadratic formula if the equation cant be factored out...

OpenStudy (anonymous):

Quadratic formula makes sense when you cannot find the factors but it is too lengthy for small questions like this

OpenStudy (anonymous):

hahaha ok thats great, i only want you to learn the process of solving thats why i ask you to use the formula...

OpenStudy (anonymous):

i will :)

OpenStudy (anonymous):

ok great...practice makes perfect..keep on practicing and solving....

OpenStudy (anonymous):

Nikko, u hv the roots of the two equations. What abt the other questions asked?? You know their answers??